OpenStudy (idealist10):

Solve the initial-value problem (x^2-1)y'-2xy=x(x^2-1), y(0)=4.

OpenStudy (idealist10):

I don't know why the answer is y=(x^2-1)[(1/2)ln abs(x^2-1)-4] when I got the answer y=x^4/(4(x^2-1))-x^2/(2(x^2-1))-4/(x^2-1).

OpenStudy (idealist10):

Here's the work: y'-2x/(x^2-1)y=x u=e^(ln abs(x^2-1))=x^2-1 (x^2-1)y'-2xy=x(x^2-1) (x^2-1)y=x^4/4-x^2/2+C y=x^4/(4(x^2-1))-x^2/(2(x^2-1))+C/(x^2-1) y(0)=-C=4 C=-4 y=x^4/(4(x^2-1))-x^2/(2(x^2-1))-4/(x^2-1)

OpenStudy (idealist10):

@goformit100 @paki @SithsAndGiggles @charlotte123

OpenStudy (anonymous):

\[\begin{align*}(x^2-1)y'-2xy&=x(x^2-1)\\ y'-\frac{2x}{x^2-1}y&=x \end{align*}\] The integrating factor is \[\begin{align*}\ln\mu(x)&=-\int\frac{2x}{x^2-1}~dx\\ \ln\mu(x)&=-\ln|x^2-1|\\ \mu(x)&=\frac{1}{x^2-1} \end{align*}\]

OpenStudy (idealist10):

I think I forgot the negative sign. So the integrating factor u=e^-ln abs(x^2-1), not e^ln abs(x^2-1), right?

OpenStudy (anonymous):

Yes, then log properties say the integrating factor is the reciprocal of what you previously got.

OpenStudy (idealist10):

Thank you. I must be careful later.

OpenStudy (anonymous):

You're welcome!